OpenStudy (anonymous):

csc 5pi/6 =? and cot 5pi/6 =? i need the answer quickly please

terenzreignz (terenzreignz):

Do you know the sine and cosine of 5pi/6?

OpenStudy (anonymous):

No the question is just like that

terenzreignz (terenzreignz):

I know, but as a reference, it pays to know (by heart) the sines and cosines of such special angles... and \(\Large \frac{5\pi}{6}\) is one such special angle... That said, what is its sine and cosine?

OpenStudy (anonymous):

I realy don't know i have an online exam now and i have to get a good score in it or i will lost 1 year in the universty )):

terenzreignz (terenzreignz):

Well then, brush up ^_^ Come, tell me what the sine and cosine of \(\Large \frac{5\pi}{6}\) are... look them up in a unit circle if you have to ...

OpenStudy (anonymous):

How can i fined it?

terenzreignz (terenzreignz):

Here... http://upload.wikimedia.org/wikipedia/commons/thumb/4/4c/Unit_circle_angles_color.svg/300px-Unit_circle_angles_color.svg.png The unit circle, just look for \(\Large \frac{5\pi }{6}\) and the left coordinate corresponds to the cosine of the angle while the right-coordinate corresponds to the sine of the angle.

OpenStudy (anonymous):

The cosin of 5pi/6 is 150 degree i think.

terenzreignz (terenzreignz):

Wrong... that's the same \(\Large \frac{5\pi}6\) only in degrees... look again (do you not know how to refer to the unit circle? D: )

OpenStudy (anonymous):

Actually no i didn't study math from 1 and half year i forgot lot of things

terenzreignz (terenzreignz):

Then... I'd suggest a comprehensive review ^_^

OpenStudy (anonymous):

And i'm not a native speaker there is lot of new words i can't understand now

terenzreignz (terenzreignz):

Where are you from, then? (Out of interest) And I'm already telling you how to use the unit circle... for each angle (including \(\Large \frac{5\pi}6\) ) there is a pair of coordinates, something like (x,y) the left-coordinate (x) is the cosine of the angle while the right coordinate (y) is the sine of the angle. I know you can do this ^_^

OpenStudy (anonymous):

I'm from Oman ^_^. i will try it

OpenStudy (anonymous):

Ahaa so -sqrt(3)/2 is cosin and 1/2 is sine right?

terenzreignz (terenzreignz):

See? You can do it ^_^ Now, the cosecant of the angle [csc] is simply the reciprocal of the sine. That means, just flip the sine (switch numerator and denominator).

OpenStudy (anonymous):

so the answer is 2/1?

terenzreignz (terenzreignz):

2. 2/1 is just 2. That's the cosecant.

OpenStudy (anonymous):

Ahaa that's really easy thanks man. i will try to solve the other one

terenzreignz (terenzreignz):

Remember...the cotangent [cot] of an angle is simply cos/sin. Divide the cosine by the sine.

OpenStudy (anonymous):

So it will be like that -sqrt(3)/2/2/1?

terenzreignz (terenzreignz):

Yes... but simplify that :)

OpenStudy (anonymous):

is - 1.73 the right answer?

terenzreignz (terenzreignz):

Oh, please don't use a calculator. Simplify by cancelling out, but leave the square roots as is.

terenzreignz (terenzreignz):

Unless it is specifically instructed to use a calculator, of course.

OpenStudy (anonymous):

so the answer will be -sqrt(3)

terenzreignz (terenzreignz):

That is correct ^_^

OpenStudy (anonymous):

Thanks a lot. By the way where are you from?

OpenStudy (anonymous):

Muslim?

terenzreignz (terenzreignz):

No ^_^ Sorry.

OpenStudy (anonymous):

Hhh that's okay bro.

OpenStudy (anonymous):

Hey do you have free time to help me with the exam i have 10 questions left

terenzreignz (terenzreignz):

But that might be considered cheating. D:

OpenStudy (anonymous):

Yaah i know but all my friends did that because we didn't stady math for like 2 years i will review it of course but i have to complete this test by tomorrow

terenzreignz (terenzreignz):

hang on...